OpenStudy (anonymous):

solve each triangle . round your answers to the nearest 10th. in ABCtriangle,a=14cm, b=9cm, c=6cm in QRPtriangle, q=12in, p=28in, r=18in inDEFtriangle, e=16yd, d=12yd, f=17yd in QRPtriagle, p=28km, q=17km, r=15km

OpenStudy (mathteacher1729):

I guess they want you to find all the angles, right?

OpenStudy (anonymous):

yess

OpenStudy (mathteacher1729):

Ok, so there is a step-by-step procedure for this given here: http://en.wikipedia.org/wiki/Solution_of_triangles#Three_sides_given_.28SSS.29 You will have to use the law of cosines (the formula given will give you the angle C). THEN use the law of sines to find the angles A and B. FIRST STEP -- just solve for c. Lemmie know what you get. :)

OpenStudy (anonymous):

ok, but r we doing the first problems because my problem is drawing it

OpenStudy (mathteacher1729):

Oh, you're supposed to draw it and then get your answers from the drawing? (like, with a ruler, compass, protractor?)

OpenStudy (anonymous):

yea like im supposed to draw a triangle thats what i was confused at

OpenStudy (mathteacher1729):

Ahh, ok. What tools do you have to draw them with? (are you using geogebra, geometer's sketchpad, or other software, or are you drawing them by hand with compass, ruler, protractor, etc?)

OpenStudy (anonymous):

it doesnt matter its freehand

OpenStudy (mathteacher1729):

Ok, well, I mean, if you want an accurate picture you should at least use a ruler and compass -- or use geogebra to generate a nice picture. Otherwise just draw any old triangle with three sides of different lengths and make sure the longest side is labeled as such. :-p

OpenStudy (anonymous):

okay so it doesnt matter how i label it

OpenStudy (anonymous):

so how do i find the angles

OpenStudy (anonymous):

i get stuck on finding the angles because i know how to do it but i get to a point where i stop

OpenStudy (mathteacher1729):

It would be kinda silly to label the smallest side with the longest length... at least try to get that accurate. :-p So for... ABC triangle,a=14cm, b=9cm, c=6cm DRAWING STEPS 1) Draw a triangle with three different side lengths. 2) Label the LONGEST side as a = 14cm 3) Label the SMALLEST side as c = 6cm. FINDING THE ANGLE STEPS To find angle C (the angle opposite the side labeled c) fill in the blanks... c^2 = a^2+b^2-2*a*b*cos(C) QUESTION ==> Do you know how to use arccos ? (inverse cosine)?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

im stuck

OpenStudy (mathteacher1729):

Ok. So let's take a step back then. Say you have 1 = sin(x) And you wanna find what x is... but you can't like "divide by sin" that doesn't make sense. You have to "UNDO" whatever the function "sin' DID to the angle x to find the result 1... so arcsin(1) = x Try it. Input arcsin(1) ... or it might be written as .... sin^(-1)(x) and see what comes out with your calc.

OpenStudy (anonymous):

okay but before i do that for this particular problem would it be sin of x/14=sin9/31

OpenStudy (mathteacher1729):

What were the steps leading up to that? Also, do you mean sin(9) divided by 31 or sin(9/31) ? Where did 31 come from?

OpenStudy (anonymous):

no i was doing an example because i need to fill in the numbers in order to find the angle but imconfused on that